# It Cannot Be Right If It Was Written by AI: On Lawyers’ Preferences of Documents Perceived as Authored by an LLM vs a Human

### v1

@@ 2.3 Use of LLMs in Legal Domain
Legal Reasoning The survey by Huang and Chang [37] notes the significant progress in NLP caused by the introduction of LLMs. However, the authors point out that the extent to which LLMs can properly reason remains unclear. Blair-Stanek et al [6] evaluated GPT-3 on the statutory-reasoning dataset SARA. The authors noted apparent limitations due to imperfect prior knowledge of U.S. statutes and poor performance on synthetic statutes not encountered during training. Studies by Nguyen et al [70, 72, 73] reported further limitations of using LLMs for legal reasoning. Additionally, Nguyen et al [71] analysed the performance of GPT-3.5 and GPT-4 on the COLIEE Task 4 dataset. The authors raised concerns about the GPT models’ ability to generalize and learn adaptable rules for unknown cases. During the preliminary exploration of the ability of GPT-3.5 for reasoning about the FOIA requests, Baron et al [3] reported ChatGPT performing below the level of an experienced FOIA reviewer. On the other hand, ChatGPT exhibited the ability to bring valuable recommendations accompanied by legal reasoning. Yu et al [100] suggested that the reasoning capabilities of LLMs can be significantly improved by Chain-of-Thought prompting and fine-tuning with explanations. The research indicates that the best results are achieved using prompts directly derived from specific legal reasoning techniques (e.g., IRAC). Guha et al [31] prepared a robust typology for organizing legal tasks and evaluating 20 LLMs from 11 different families to provide a benchmark for the legal reasoning capabilities of LLMs. Following the modified IRAC structure, the evaluation revealed diverging performance levels, with GPT-4 emerging as the most successful model. Generally, LLMs perform better on classification tasks than those focused on application. Kang et al [46] evaluated GPT-3.5’s ability to conduct IRAC analysis. The research found that powerful LLMs can provide reasonable answers but mostly fail to yield correct reasoning paths. Finally, Janatian et al [41] suggest using GPT-4 to extract pathways from real-world legislation to support the development of legal expert systems. Their evaluation yielded that 60% of the generated pathways were equivalent or superior to manually created ones.
@@
Support for Legal Research and eDiscovery Various tasks related to legal research, legal information retrieval or eDiscovery can be supported by LLMs. Savelka et al [83] demonstrated the effectiveness of GPT-4 augmented with a legal information retrieval module, which significantly improved the accuracy of explanations of legal concepts. These findings correspond with Blair-Stanek et al [7], who used non-augmented models. The authors reported poor performance of most state-of-the-art LLMs (including GPT-4) in basic legal text-handling tasks, referring to them as ’sloppy paralegals’. Integrating LLMs with knowledge bases [19] and other tools allowing, e.g., factual lookups [84] and gathering references [66] can boost performance. However, the existing research is focused on domains outside of law. Huang et al [38] emphasized the importance of domain-specific knowledge over the general experiences distilled from ChatGPT, as demonstrated through supervised fine-tuning tasks. Henseler and van Beek [35] used GPT-4 as a supportive tool for handling data in digital investigation, entrusting it with tasks like summarization, evaluation and visualization of chat messages. Their work showcased the potential of LLMs in eDiscovery processes. Ioannidis et al [39] utilized LLMs for regulatory compliance, providing identification, summarization and the impact level assessment of legal rules impacting businesses. Lam et al [54] employed LLMs to streamline contract drafting processes, facilitating tasks such as template selection or clause modification to suit specific contexts agreed upon by parties.
@@
Access to Justice/Law Pre-trained language models offer a significant potential for increasing the accessibility of domain-specific knowledge to laypeople. Classical information retrieval can be paired with pre-trained models to substantially improve domain expert advice. Metzler et al [64] introduced a consolidated model combining multi-task learning and zero- or few-shot learning, which may serve as a foundation for including ChatGPT into various pipelines for improved question-answering capabilities. Addressing the procedural and financial barriers to legal services, LLMs have the potential to improve access to justice and government services [9]. Research has explored various avenues for utilizing LLMs, including support to lobbying activities [68], mediation support [97], and mapping laypeople’s fact-based narrative to legal issues [96]. Chien et al [17] highlighted the deployment of GPT-powered chatbots in the judiciary, improving efficiency in addressing the legal needs of low-income groups. Goodson and Lu [27] reported a proof-of-concept utilization of LLMs to streamline the legal aid intake process and support the initial legal triage. Such use addresses a known bottleneck in legal aid delivery. Laypeople often utilize LLM-based tools for self-help in legal issues. They mainly seek statements about applicable laws, procedural guidance and directories of potential services or webpages delivering further assistance [32].
@@
Legal Judgment Prediction Legal judgment prediction aims to predict judgment results automatically. Experiments showcase the potential of LLMs for this task. Hamilton [33] demonstrated GPT-2’s ability to achieve better-than-random accuracy in simulating judicial rulings of the US Supreme Court from 2010 to 2016. Similarly, zero-shot prediction on the Chinese criminal case dataset using GPT-3 reported by Jiang and Yang [42] outperformed the baseline. The investigation by Trautmann et al [91] focused on the prediction of judgments of the European Court of Human Rights and the Federal Supreme Court of Switzerland. The study found that the LLMs exhibited superior zero-shot predictive capabilities compared to baseline models. However, the reported results fell short of existing supervised approaches. Wu et al [98] demonstrated the possibility for improved performance by combining LLMs with domain models. The introduction of GPT-4 presented a significant development, yielding remarkably superior prediction results compared to previous models and existing approaches [88].
@@
Annotation LLMs can automatically annotate textual elements, significantly reducing effort and costs associated with empirical analysis and legal review. Drapal et al [24] proposed using GPT-4 to support thematic analysis in empirical legal studies with promising results in zero-shot classification and the potential for further improvement with expert feedback. Savelka [80] and Savelka and Ashley [81] evaluated GPT-3.5’s effectiveness in contract review, statutory and regulatory provisions investigation, and case-law analysis. Gray et al [29] utilized GPT-4 and GPT-4-Turbo to create sentence-level annotations for factor analysis. Savelka et al [82] found GPT-4 comparable to well-trained student annotators in analyzing textual data. Oliveira et al [75] explored the use of GPT-3 in a pipeline integrating LLMs and weak supervision for named entity recognition tasks. They observed comparable performance to traditional manual text labelling methods at a lower cost.
@@ 3 Evaluation Experiment
We conducted an evaluation experiment of two documents (Brief and Verbose) detailed in Subsection 3.1 and two groups of evaluators described in Subsection 3.2. Group A received the Brief document labelled as AI-generated and the Verbose one labelled Human-crafted. Group B received the documents with opposite labelling. Participants were instructed to evaluate the language quality and correctness of the documents through a survey. They were allowed to provide open-ended comments to their evaluation. Details about the survey can be found in Subsection 3.3.
@@ 3.1 Documents
We prepared two written acknowledgements of debt. Both documents were drafted to fit on one page. One document was prepared to be as brief as possible (Brief), while the other was drafted to contain some additional information (Verbose). Both documents were prepared to comply with standard practices (e.g., designation of parties, legal jargon used).
@@
The Verbose document contained the structured designation of the debtor and the creditor (name, surname, date of birth, address) followed by the headline ’Acknowledgment of Debt’. The document contained an explicit acknowledgement of debt required by law. Subsequently, the document outlined the origin of the debt (failure to pay the lease) in a structured and more detailed manner. Finally, the document contained the due date and the confirmation that it was not written under duress. The verbose document was approximately 250 words in length.
@@
Further, each document was modified into two versions. One version was labelled ’AI-GENERATED DOCUMENT’ in both the header and the footer in blue highlight. The other version was labelled ’HUMAN-CRAFTED DOCUMENT’ in both the header and the footer in yellow highlight. We avoided using green and red highlights for their well-known associations to correct and incorrect options.
@@
Figure 2: The figure contains snippets of two variants of the Verbose document. One is designated as ’AI-GENERATED DOCUMENT’ (left) in its header, and the other as ’HUMAN-CRAFTED DOCUMENT’ (right). The same designation appears also in the footer of every document.
@@ 3.2 Participants
We recruited participants through a mailing campaign at an R1 university in Europe, and calls for participants distributed via Facebook and Twitter, reaching approximately 1,200 people. A total of 89 prospective subjects expressed their interest in participating. Of the 89 participants, 53 were law students enrolled in a law degree granting program, and 36 were lawyers. We used stratified random sampling to divide participants into two groups: Group A and Group B.
@@
Originally, we planned to discard the answers from participants who completed the survey in under ten minutes. We did not discard any answers for this reason as all of the participants took longer to finish the survey.
@@ 3.3 Survey
The evaluation was conducted online via Microsoft Forms. Group A was presented with the Brief document labelled as AI-generated and the Verbose document labelled as Human-crafted. Group B was presented with the Brief document labelled as Human-crafted and the Verbose document labelled as AI-generated. Participants were not made aware that both documents were Human-crafted. Survey was distributed to all the participants on January 30, 2024. Participants were requested to finish their evaluation by February 11, 2024.
@@
    Correctness: The degree to which the document is correct. The participants were tasked to consider the fulfilment of legally required conditions and factual and formal coherence.
Participants were also asked to provide short explanations (100 words suggested) for each scoring question. Finally, the evaluation included an open-ended question to collect participants’ opinions on the possibility of generating debt acknowledgement automatically and achieving output quality comparable to humans.
@@
Participants were required to answer all four scoring questions (Questions 1, 3, 5, and 7) before the survey could be submitted. Answering open-ended questions (Questions 2, 4, 6, 8, and 9) was not mandatory.
The evaluation of the survey was done at the level of scoring and open-ended questions. The scoring questions were evaluated quantitatively, and in particular in terms of average scores and in terms of preference for human-crafted or AI-generated documents with respect to the brief or verbose nature of the documents.
@@ 4.1 Document Preference
While both documents were authored to be of comparable quality it appears that the verbose one was preferred by the participants, especially in terms of correctness. The mean overall correctness rating of the brief document was 4.24 (39$\times$ presented as AI-generated and 36$\times$ as human-authored) compared to the 4.67 of the verbose one (36$\times$ AI and 39$\times$ human). Seen side-by-side, the participants judged the verbose document as more correct than the brief one 34 times (out of 75). The brief document was deemed as more correct only 9 times. In the remaining 32 times the documents were evaluated as comparable. In terms of language, the trend was still there (4.39 verbose versus 4.13 brief) where the verbose document was preferred 26 times compared to 12 preferences for the brief one (37 neither). Figure 3 provides additional detail into how the two documents were rated by the participants. While the comparison of the documents is not the subject of this study it presents an important aspect that needs to be considered when interpreting the results below.
Figure 3: The figure summarizes the participants’ preferences between the two documents (Brief and Verbose). The top two charts show the distribution of scores awarded to each document in terms of their Correctness and Language (1–worst; 5–best). The bottom two charts present the results of the side-by side comparisons, showing how many times each of the documents was preferred (if any). Overall, a clear preference for the Verbose document over the Brief one can be observed.
@@ 4.2 Correctness
Figure 4 shows the overall evaluation of the documents in terms of correctness, focusing on whether a document was labeled as human-crafted or AI-generated. It appears that the participants clearly preferred a document when it was presented as human-crafted over a document labeled as AI-generated. Specifically, the mean evaluation of the two documents when labeled as human-crafted was 4.69 as compared to 4.21 when the same documents were marked as AI-generated. The side-by-side comparison presents a clear message when the human-crafted designation yielded 35 preferences over the mere 7 of the documents perceived as generated by AI (33 neither). This difference is statistically significant by Fisher exact test ($p<10^{-5}$).
@@
Figure 4: The figure summarizes the participants’ preferences between the documents when labeled as AI-generated versus human-crafted in terms of their Correctness. The top two charts show the distribution of scores awarded to the documents carrying the “AI“ (green) or “human“ (blue) labels (1–worst; 5–best). The bottom chart presents the results of the side-by side comparison, showing how many times each of the labels was preferred (if any). Overall, a clear preference for the documents labeled as human-crafted over those labeled as AI-generated can be observed.
@@
Figure 5 provides insight into the interaction between the AI/human-authored designation and the individual documents. While the verbose document appears to be clearly preferred by the participants overall (Figure 3), the effect is negated if it is marked as AI-generated, and the brief document as authored by a human. In that case, we can even observe that the brief document was deemed as slightly more correct on average (4.5 brief versus 4.44 verbose). In terms of the side-by-side comparisons, the verbose document (labeled as AI-generated) was preferred 7 times whereas the one marked as authored by a human was deemed more correct 9 times (20 neither). The overall preference for the verbose document was greatly boosted when this one was labeled as produced by a human. Specifically, the mean correctness rating of the verbose document marked as coming from a human was 4.87 compared to the mere 4.0 of the brief document presented as AI-generated. The verbose document was preferred in 26 out of 39 instances while the brief document was not preferred a single time (i.e., 13 neither).
@@
Figure 5: The figure summarizes the participants’ preferences between the documents when labeled as AI-generated versus human-crafted in terms of their Correctness when the document is taken into account (Brief and Verbose). The top two charts show the distribution of scores awarded to the documents carrying the “AI“ (green) or “human“ (blue) labels (1–worst; 5–best). The bottom charts present the results of the side-by side comparisons, showing how many times each of the labels was preferred (if any) per document. While the Verbose document is clearly preferred overall, the AI-generated label appears to largely mitigate the effect in case of when attached to the Verbose document and largely amplify it when put on the Brief document.
@@
Figure 6 focuses on the interaction between the participants’ background (i.e., law student versus lawyer) and the AI/human designation of a document. It appears that law students were heavily influenced by the designation, preferring a document marked as human-authored (4.81 labeled as human versus 4.19 for AI). Specifically, the law students preferred the document marked as coming from a human in 21 instances compared to only 2 preferences for a document presented as generated by AI (19 neither). Lawyers were less susceptible to the AI/human designation while still preferring a document when perceived as coming from a human (4.55 human vs 4.24 AI). They expressed preference for a human-authored document in 14 cases compared to 5 instances when a document marked as AI-generated was preferred (14 neither). Note that the result is not statistically significant by Fisher exact test ($p=0.2221$).
@@
Figure 6: The figure summarizes the participants’ preferences between the documents when labeled as AI-generated versus human-crafted in terms of their Correctness when the participants’ background is taken into account (a law student and a lawyer). The top two charts show the distribution of scores awarded to the documents carrying the “AI“ (green) or “human“ (blue) labels (1–worst; 5–best). The bottom charts present the results of the side-by side comparisons, showing how many times each of the labels was preferred (if any) per participants’ background. It appears the AI-generated label has much larger effect on the law students as compared to the lawyers.
@@
  • Wording and conventions – whether the document follows common legal conventions and legal terms.
All the themes were either mentioned in a positive or a negative sense. Examples of positive mentions are, e.g., “fulfils legal requirements,” or “correctly identifies the parties, the related documents (contract) and the correct Act and Section”. Examples of negative mentions are, e.g., “document does not contain the specific provisions of the referenced contract” or “paragraphs appear inconsistent and information is not conveyed smoothly”.
@@
We see the following if we take a closer look at the individual occurrences of these thematic categories. Compliance with legal requirements is mentioned exclusively positively. This is because both documents objectively and deliberately meet the legal requirements. This topic is often mentioned in the context of the relevant Section 2053 of the Czech Civil Code, which sets out the conditions for the acknowledgement of the debt, i.e. as ”the document contains what it should contain according to the relevant provision [of Civil Code]” or ”it fulfils the conditions in Section 2053 of the Civil Code”. Often this theme is mentioned in the context of the theme of expected legal consequences, i.e. as ”the document I consider to meet the requirements of Section 2053 and I consider that the document should trigger the expected legal consequences of debt acknowledgement, as set out in the document” or ”the document meets the requirements of debt acknowledgement, i.e. it contains (i) the reason for the debt (debt from the lease agreement) and (ii) the amount owed (CZK 25,000). The document contains placeholders for identifying the debtor and the creditor, the completion of which will produce the desired legal consequences when the document is signed”.
Similarly, this theme also appears frequently in combination with the theme of formal correctness in a positive sense, i.e., as, for example, ”the debt is acknowledged by the document to the extent and in the amount required by Section 2053 of the Civil Code, and this is done with sufficient precision” or ”I consider the requirements of Section 2053 of the Civil Code to have been fulfilled. The document is materially and formally coherent.” In general, answers have shown that there is broad agreement amongst evaluators this requirement is met and accordingly this theme is consistently mentioned.
The theme of coherence and logical structure contains negative comments on redundancy, e.g. ”[…] the terminology is followed, although some of the wording seems redundant”, or i.e. comments on reduced clarity, e.g. ”I had to think about which rent for which month is due” or ”in the introduction of the document we learn that the debt was incurred as rent due on 15th January 2015 and then we learn that the debtor did not pay the rent until 2023, the document then appears less consistent”. In a positive sense, the document is commented on as ”it is written in a very clear and well-organised manner” or ”the document proceeds logically and clearly, emphasising the important facts (reason and amount of the debt), […]”.

### v2

@@ 2.3 Use of LLMs in Legal Domain
Legal Reasoning The survey by Huang and Chang [39] notes the significant progress in NLP caused by the introduction of LLMs. However, the authors point out that the extent to which LLMs can properly reason remains unclear. Blair-Stanek et al [6] evaluated GPT-3 on the statutory-reasoning dataset SARA. The authors noted apparent limitations due to imperfect prior knowledge of U.S. statutes and poor performance on synthetic statutes not encountered during training. Studies by Nguyen et al [73, 75, 76] reported further limitations of using LLMs for legal reasoning. Additionally, Nguyen et al [74] analysed the performance of GPT-3.5 and GPT-4 on the COLIEE Task 4 dataset. The authors raised concerns about the GPT models’ ability to generalize and learn adaptable rules for unknown cases. During the preliminary exploration of the ability of GPT-3.5 for reasoning about the FOIA requests, Baron et al [3] reported ChatGPT performing below the level of an experienced FOIA reviewer. On the other hand, ChatGPT exhibited the ability to bring valuable recommendations accompanied by legal reasoning. Yu et al [104] suggested that the reasoning capabilities of LLMs can be significantly improved by Chain-of-Thought prompting and fine-tuning with explanations. The research indicates that the best results are achieved using prompts directly derived from specific legal reasoning techniques (e.g., IRAC). Guha et al [33] prepared a robust typology for organizing legal tasks and evaluated 20 LLMs from 11 different families to provide a benchmark for the legal reasoning capabilities of LLMs. Following the modified IRAC structure, the evaluation revealed diverging performance levels, with GPT-4 emerging as the most successful model. Generally, LLMs perform better on classification tasks than those focused on application. Kang et al [48] evaluated GPT-3.5’s ability to conduct IRAC analysis. The research found that powerful LLMs can provide reasonable answers but mostly fail to yield correct reasoning paths. Finally, Janatian et al [43] suggest using GPT-4 to extract pathways from real-world legislation to support the development of legal expert systems. Their evaluation yielded that 60% of the generated pathways were equivalent or superior to manually created ones.
@@
Support for Legal Research and eDiscovery Various tasks related to legal research, legal information retrieval, drafting or eDiscovery can be supported by LLMs. Savelka et al [86] demonstrated the effectiveness of GPT-4 augmented with a legal information retrieval module, which significantly improved the accuracy of explanations of legal concepts. These findings correspond with Blair-Stanek et al [7], who used non-augmented models. The authors reported poor performance of most state-of-the-art LLMs (including GPT-4) in basic legal text-handling tasks, referring to them as ’sloppy paralegals’. Integrating LLMs with knowledge bases [20] and other tools allowing, e.g., factual lookups [87] and gathering references [69] can boost performance. However, the existing research is focused on domains outside of law. Huang et al [40] emphasized the importance of domain-specific knowledge over the general experiences distilled from ChatGPT, as demonstrated through supervised fine-tuning tasks. Henseler and van Beek [37] used GPT-4 as a supportive tool for handling data in digital investigation, entrusting it with tasks like summarization, evaluation and visualization of chat messages. Their work showcased the potential of LLMs in eDiscovery processes. Ioannidis et al [41] utilized LLMs for regulatory compliance, providing identification, summarization and the impact level assessment of legal rules impacting businesses. Lam et al [56] employed LLMs to streamline contract drafting processes, facilitating tasks such as template selection or clause modification to suit specific contexts agreed upon by parties. Similarly, Trozze et al [95] used LLMs to support legal drafting in securities cases, noting models’ inability to reason but leading to reduced drafting time. In the first randomized controlled trial of the effect of AI assistance, Choi et al [19] found slight and inconsistent improvement in the quality of legal analysis but significant and consistent increases in speed.
@@
Access to Justice/Law Pre-trained language models offer a significant potential for increasing the accessibility of domain-specific knowledge to laypeople. Classical information retrieval can be paired with pre-trained models to improve domain expert advice substantially. Metzler et al [67] introduced a consolidated model combining multi-task learning and zero- or few-shot learning, which may serve as a foundation for including ChatGPT into various pipelines for improved question-answering capabilities. Addressing the procedural and financial barriers to legal services, LLMs can potentially improve access to justice and government services [9]. Research has explored various avenues for utilizing LLMs, including support to lobbying activities [71], mediation support [101], and mapping laypeople’s fact-based narrative to legal issues [100]. Chien et al [17] highlighted the deployment of GPT-powered chatbots in the judiciary, improving efficiency in addressing the legal needs of low-income groups. Goodson and Lu [29] reported a proof-of-concept utilization of LLMs to streamline the legal aid intake process and support the initial legal triage. Such use addresses a known bottleneck in legal aid delivery. Laypeople often utilize LLM-based tools for self-help in legal issues. They mainly seek statements about applicable laws, procedural guidance and directories of potential services or webpages delivering further assistance [34].
@@
Legal Judgment Prediction Legal judgment prediction aims to predict judgment results automatically. Experiments showcase the potential of LLMs for this task. Hamilton [35] demonstrated GPT-2’s ability to achieve better-than-random accuracy in simulating judicial rulings of the U.S. Supreme Court from 2010 to 2016. Similarly, zero-shot prediction on the Chinese criminal case dataset using GPT-3 reported by Jiang and Yang [44] outperformed the baseline. The investigation by Trautmann et al [94] focused on the prediction of judgments of the European Court of Human Rights and the Federal Supreme Court of Switzerland. The study found that the LLMs exhibited superior zero-shot predictive capabilities compared to baseline models. However, the reported results fell short of existing supervised approaches. Wu et al [102] demonstrated the possibility for improved performance by combining LLMs with domain models. The introduction of GPT-4 presented a significant development, yielding remarkably superior prediction results compared to previous models and existing approaches [91].
@@
Annotation LLMs can automatically annotate textual elements, significantly reducing the effort and costs associated with empirical analysis and legal review. Drapal et al [26] proposed using GPT-4 to support thematic analysis in empirical legal studies with promising results in zero-shot classification and the potential for further improvement with expert feedback. Savelka [83] and Savelka and Ashley [84] evaluated GPT-3.5’s effectiveness in contract review, statutory and regulatory provisions investigation, and case-law analysis. Gray et al [31] utilized GPT-4 and GPT-4-Turbo to create sentence-level annotations for factor analysis. Savelka et al [85] found GPT-4 comparable to well-trained student annotators in analyzing textual data. Oliveira et al [78] explored the use of GPT-3 in a pipeline integrating LLMs and weak supervision for named entity recognition tasks. They observed comparable performance to traditional manual text labelling methods at a lower cost.
@@ 3 Evaluation Experiment
We conducted an evaluation experiment of two human-crafted documents (Brief and Verbose) detailed in Subsection 3.1 and two groups of evaluators described in Subsection 3.2. Group A received the Brief document labeled as AI-generated and the Verbose one labeled human-crafted. Group B received the documents with opposite labelling. Participants were instructed to evaluate the language quality and correctness of the documents through a survey. They were allowed to provide open-ended comments to their evaluation. Details about the survey can be found in Subsection 3.3.
@@ 3.1 Documents
We prepared two written acknowledgements of debt. Both documents were human-crafted without the use of LLMs and drafted to fit on one page. One document was prepared to be as brief as possible (Brief), while the other was drafted to contain some additional information (Verbose). Both documents were prepared to comply with standard practices (e.g., designation of parties, legal jargon used).
@@
The Verbose document contained the structured designation of the debtor and the creditor (name, surname, date of birth, address) followed by the headline ’Acknowledgement of Debt’. The document contained an explicit acknowledgement of debt required by law. Subsequently, the document outlined the origin of the debt (failure to pay the lease) in a structured and more detailed manner. Finally, the document contained the due date and the confirmation that it was not written under duress. The Verbose document was approximately 250 words in length.
@@
Further, each document was modified into two versions. One version was labeled ’AI-GENERATED DOCUMENT’ in both the header and the footer in blue highlight. The other version was labeled ’HUMAN-CRAFTED DOCUMENT’ in both the header and the footer in yellow highlight. We avoided using green and red highlights for their well-known associations to correct and incorrect options.
@@
Figure 2: The figure contains snippets of two variants of the Verbose document. One is designated as ’AI-GENERATED DOCUMENT’ (left) in its header, and the other as ’HUMAN-CRAFTED DOCUMENT’ (right). The same designation appears also in the footer of every variant.
@@ 3.2 Participants
We recruited participants through a mailing campaign at Masaryk University in Czechia and calls for participants distributed via Facebook and X, reaching approximately 1,200 people.
Prior to taking the survey, prospective participants were informed about the scope of their participation, data usage, and rights, including the right to withdraw at any time. The data was collected anonymously, and no personally identifiable information was shared with the research team. No piece of collected data can be tied back to a specific individual.
A total of 89 prospective participants expressed their interest in participating. Of the 89 participants, 53 were law students enrolled in a law degree-granting program, and 36 were lawyers. We used stratified random sampling to divide participants into Group A and Group B.
@@
Initially, we planned to discard the answers from participants who completed the survey in under ten minutes. However, we did not discard any answers for this reason, as all of the participants took longer to finish the survey.
@@ 3.3 Survey
The survey was conducted online via Microsoft Forms. Group A was presented with the Brief document labeled as AI-generated and the Verbose document labeled as human-crafted. Group B was presented with the Brief document labeled as human-crafted and the Verbose document labeled as AI-generated. Participants were unaware that both documents were human-crafted and that no AI-based tools were used. The survey was distributed to all the participants on January 30, 2024. Participants were requested to finish their evaluation by February 11, 2024.
@@
    Correctness: The degree to which the document is correct. The participants were to evaluate the fulfilment of legally required conditions and factual and formal coherence.
Participants were also asked to provide brief explanations (suggested length of up to 100 words) for each score given. Finally, the evaluation included an open-ended question to collect participants’ opinions on the feasibility of generating debt acknowledgements automatically and achieving output quality comparable to that of humans.
@@
Before the survey could be submitted, participants were required to answer all four scoring questions (Questions 1, 3, 5, and 7). Answering open-ended questions (Questions 2, 4, 6, 8, and 9) was not mandatory.
We evaluated the survey using both quantitative and qualitative methods. Quantitatively, we analyzed participants’ scores, focusing on average scores and preferences for documents labeled as human-crafted or AI-generated, with a further distinction between the Brief and the Verbose documents. Qualitatively, we examined open-ended responses.
@@ 4.1 Document Preference
While both documents were authored to be of comparable quality, it appears that the participants preferred the Verbose one, especially in terms of correctness. The mean overall correctness rating of the Brief document was 4.24 (39$\times$ presented as AI-generated and 36$\times$ as human-crafted) compared to the 4.67 of the Verbose one (36$\times$ AI and 39$\times$ human). Seen side-by-side, the participants judged the Verbose document as more correct than the Brief one 34 times (out of 75). The Brief document was deemed as more correct only 9 times. In the remaining 32 times the documents were evaluated as comparable. In terms of language quality, the trend was still there (4.39 Verbose versus 4.13 Brief), where the Verbose document was preferred 26 times compared to 12 preferences for the Brief one (37 neither). Figure 3 provides additional detail into how the participants rated the two documents. While the comparison of the documents is not the subject of this study, it presents an important aspect that needs to be considered when interpreting the results below.
Figure 3: The figure summarizes the participants’ preferences between the two documents (Brief and Verbose). The top two charts show the distribution of scores awarded to each document in terms of their Correctness and Language Quality (1–worst; 5–best). The bottom two charts present the results of the side-by-side comparisons, showing how many times each of the documents was preferred (if any). Overall, a clear preference for the Verbose document over the Brief one can be observed.
@@ 4.2 Correctness
Figure 4 shows the overall evaluation of the documents in terms of correctness, focusing on whether a document was labeled as human-crafted or AI-generated. The participants preferred a document when it was presented as human-crafted over a document labeled as AI-generated. Specifically, the mean evaluation of the two documents when labeled as human-crafted was 4.69, compared to 4.21 when the same documents were marked as AI-generated. The side-by-side comparison presents a clear message when the human-crafted designation yielded 35 preferences over the mere 7 of the documents perceived as generated by AI (33 neither). This difference is statistically significant by Fisher exact test ($p<10^{-4}$ using Bonferroni correction for 4 comparisons).
@@
Figure 4: The figure summarizes the participants’ preferences between the documents when labeled as AI-generated versus human-crafted in terms of their Correctness. The top two charts show the distribution of scores awarded to the documents carrying the “AI“ (green) or “human“ (blue) labels (1–worst; 5–best). The bottom chart presents the results of the side-by-side comparison, showing how many times each of the labels was preferred (if any). Overall, a clear preference for the documents labeled as human-crafted over those labeled as AI-generated can be observed.
@@
Figure 5 provides additional insight into the preferences, focusing on the interaction between the AI-authored and human-crafted designation and the individual documents. While the Verbose document appears to be preferred by the participants overall (Figure 3), the effect is negated if it is marked as AI-generated and the Brief document as authored by a human. In that case, we can even observe that the Brief document was deemed slightly more correct on average (4.5 Brief versus 4.44 Verbose). In terms of the side-by-side comparisons, the Verbose document (labeled as AI-generated) was preferred 7 times, whereas the one marked as authored by a human was deemed more correct 9 times (20 neither). The overall preference for the Verbose document was significantly boosted when this one was labeled as produced by a human. Specifically, the mean correctness rating of the Verbose document marked as coming from a human was 4.87 compared to the mere 4.0 of the Brief document presented as AI-generated. The Verbose document was preferred in 26 out of 39 instances, while the Brief document was not preferred a single time (i.e., 13 neither was preferred).
@@
Figure 5: The figure summarizes the participants’ preferences between the documents when labeled as AI-generated versus human-crafted in terms of their Correctness when the document is taken into account (Brief and Verbose). The top two charts show the distribution of scores awarded to the documents carrying the “AI“ (green) or “human“ (blue) labels (1–worst; 5–best). The bottom charts present the results of the side-by-side comparisons, showing how many times each of the labels was preferred (if any) per document. While the Verbose document is clearly preferred overall, the AI-generated label appears to largely mitigate the effect in case of when attached to the Verbose document and largely amplify it when put on the Brief document.
@@
Figure 6 focuses on the interaction between the participants’ background (i.e., law student versus lawyer) and a document’s AI/human designation. Note that the result is not statistically significant based on the Fisher exact test ($p=0.2221$). However, it appears that the designation of origin somewhat more influenced law students, preferring a document marked as human-crafted (4.81 labeled as human versus 4.19 for AI). Specifically, the law students preferred the document marked as coming from a human in 21 instances compared to only 2 preferences for a document presented as generated by AI (19 neither). Lawyers appear to have been less susceptible to the AI/human designation while still preferring a document when presented as coming from a human (4.55 human vs 4.24 AI). They preferred a human-crafted document in 14 cases compared to 5 instances when a document marked as AI-generated was preferred (14 neither).
@@
Figure 6: The figure summarizes the participants’ preferences between the documents when labeled as AI-generated versus human-crafted in terms of their Correctness when the participants’ background is taken into account (a law student and a lawyer). The top two charts show the distribution of scores awarded to the documents carrying the “AI“ (green) or “human“ (blue) labels (1–worst; 5–best). The bottom charts present the results of the side-by-side comparisons, showing how many times each of the labels was preferred (if any) per participants’ background. It appears the AI-generated label has much larger effect on the law students as compared to the lawyers.
@@
  • Wording and conventions – whether the document follows common legal conventions and terms.
All the themes were either mentioned in a positive or a negative sense. Examples of positive mentions are e.g.,
  • “Fulfils legal requirements.”
  • “Correctly identifies the parties, the related documents (contract) and the correct Act and its Section.”
Examples of negative mentions are e.g.,
  • “Document does not contain the specific provisions of the referenced contract.”
  • “Paragraphs appear inconsistent, and information is not conveyed smoothly.”
@@
Upon closer examination of the individual occurrences of these thematic categories, we see the following. Compliance with legal requirements is mentioned exclusively positively. This is because both documents objectively and deliberately meet the legal requirements. This topic is often mentioned in the context of the relevant Section 2053 of the Czech Civil Code, which sets out the conditions for the acknowledgement of the debt, e.g.,
  • ”The document contains what it should contain according to the relevant provision [of the Civil Code].”
  • ”It fulfils the conditions in Section 2053 of the Civil Code”.
Furthermore, the theme is often mentioned in the context of expected legal consequences, i.e.,
  • ”The document meets the requirements of Section 2053, and the document should trigger the expected legal consequences of debt acknowledgement.”
  • ”The document meets the requirements of debt acknowledgement, i.e. it contains (i) the reason for the debt (debt from the lease agreement) and (ii) the amount owed (CZK 25,000). The document contains placeholders for identifying the debtor and the creditor, the completion of which will produce the desired legal consequences once the document is signed.”
Similarly, the theme also appears frequently in combination with the theme of formal correctness in a positive sense. The responses showed that there is a broad consensus that formal correctness is linked to expected legal consequences:
  • ”The debt is acknowledged by the document to the extent and in the amount required by Section 2053 of the Civil Code, and this is done with sufficient precision.”
  • ”I consider the requirements of Section 2053 of the Civil Code to have been fulfilled. The document is materially and formally coherent.”
The theme of coherence and logical structure contains negative comments on redundancy or insufficient clarity of text, e.g.
  • ”[…] the correct terminology is used, although some of the wording seems redundant.”
  • ”I had to think about which rent for which month is due.”
  • ”In the document’s introduction, we learn that the debt was incurred as rent due on 15th January 2015, and then we learn that the debtor did not pay the rent until 2023; as a result, the document appears less consistent.”
In a positive sense, the document is commented on as written clearly and well, proceeding logically and clearly and emphasizing the important facts.